OpenStudy (anonymous):

WILL MEDAL AND FAN ~ Evaluate the function rule for the given value y = 4 x 2 ^x for x = -6

OpenStudy (welshfella):

= 4*-6 * 2 ^ -6 = -24 * 1/2^6 can you continue?

OpenStudy (anonymous):

I think its 1/32 not sure

OpenStudy (welshfella):

2^6 = 64 2^-6 = 1/64

OpenStudy (anonymous):

-48??

OpenStudy (welshfella):

I'm assuming the first 'x' in y = 4 x 2^x does not mean 'times' - am i correct?

OpenStudy (anonymous):

Yes "2^x" just means I'm doubling 2

OpenStudy (welshfella):

not that one - the x after the 4

OpenStudy (anonymous):

no that's multiplying

OpenStudy (welshfella):

2^x means you are taking 2 to the power x for example 2^3 means 2*2*2 = 8 - 2 multiplied by itself 3 times

OpenStudy (welshfella):

oh ok maybe u should use * for multiply to avoid confusion

OpenStudy (welshfella):

so we have